### v2

# Gelation in Vector Multiplicative Coalescence and Extinction in Multi-Type Poisson Branching Processes

Abstract.
In this note, we present a novel connection between a multi-type (vector) multiplicative coalescent process and a multi-type branching process with Poisson offspring distributions. More specifically, we show that the equations that govern the phenomenon of gelation in the vector multiplicative coalescent process are equivalent to the set of equations that yield the extinction probabilities of the corresponding multi-type Poisson branching process. We then leverage this connection with two applications, one in each direction. The first is a new quick proof of gelation in the vector multiplicative coalescent process, and the second is a new series expression for the extinction probabilities of the multi-type Poisson branching process.
Key words and phrases:
Vector-multiplicative coalescent, Gelation, Multitype branching processes, Extinction, Lambert-Euler inversion, Random graphs
1. Introduction
The main motivation for this work is paper [16] where the authors used the multidimensional Lambert-Euler inversion in the analysis of the vector multiplicative coalescent process. More specifically, they solved the multidimensional Lambert-Euler inversion and utilized the solution to investigate the gelation phenomenon in the coagulation model with the vector-multiplicative kernel. Our goal in this note is to present an interesting connection between the vector multiplicative coalescent process and a multi-type branching process with Poisson offspring distributions. As a consequence, we give a quick proof of the gelation phenomenon. Additionally, this connection leads to a series expression for the extinction probabilities of the multi-type Poisson branching process. While the correspondence highlighted here is fairly straightforward, it illuminates a deep connection between these two paradigms which we hope will continue to reveal new insights and potential for cross-disciplinary research.We begin by recalling some of the notation and concepts from the theory of vector multiplicative coalescent processes. For a detailed account of this theory, we refer the reader to [15, 16].
1.1. Vector multiplicative coalescent processes
Roughly speaking, a vector multiplicative coalescent process is a Markov process that describes cluster merger dynamics, where the weight of each cluster is a finite-dimensional vector. Following the convention in [16], to define this process formally, we use bra-ket notation. Specifically, $|{\bf x}\rangle$ will denote the column vector representation of vector ${\bf x}\in\mathbb{R}^{k}$, and $\langle{\bf x}|$ will denote the row vector representation of vector ${\bf x}\in\mathbb{R}^{k}$. For $c\in\mathbb{R}$ and ${\bf x}\in\mathbb{R}^{k}$, $\,c|{\bf x}\rangle$ will represent the product $c{\bf x}$, a column vector. Correspondingly, $\langle{\bf x}|{\bf y}\rangle=\langle{\bf y}|{\bf x}\rangle$ will be the dot product of ${\bf x}$ and ${\bf y}$ in $\mathbb{R}^{k}$. Finally, for a matrix $M\in\mathbb{R}^{k\times k}$, $\langle{\bf x}|M|{\bf y}\rangle$ will represent the product ${\bf x}^{\sf T}M{\bf y}$ resulting in a scalar.
Consider a system with $k$ types of particles: $1,\ldots,k$, or equivalently, a system where all the particles are partitioned into $k$ parts. For a given vector ${\boldsymbol{\alpha}}=(\alpha_{1},\alpha_{2},\ldots,\alpha_{k})\in(0,\infty)^{k}$, the process begins with $\langle{\boldsymbol{\alpha}}n|{\boldsymbol{1}}\rangle$ singletons distributed between the $k$ types so that for each $i$, there are $\alpha_{i}\,n$ particles of type $i$.
Let $V\in\mathbb{R}^{k\times k}$ be a nonnegative, irreducible and symmetric matrix, which we will refer to as the partition interaction matrix. A particle of type $i$ bonds with a particle of type $j$ with the intensity rate $v_{i,j}/n$, where $v_{i,j}$ is the $(i,j)$ element in the matrix $V$. The bonds are formed independently. This process is called vector-multiplicative coalescent. Note that the irreducible condition on $V$ ensures that particles of each type could ultimately coalesce with particles of every other type at some time in the vector-multiplicative coalescent process.
The size of a cluster of particles is represented by a $k$-dimensional vector ${\bf x}=(x_{1},x_{2},\ldots,x_{k})\in\mathbb{Z}_{+}^{k}$ such that $\sum_{i=1}^{k}x_{i}>0$, where each $x_{i}$ represents the number of particles of type $i$ (or in part $i$) in the cluster. Each cluster of weight ${\bf x}$ bonds together $x_{1},x_{2},\ldots,x_{k}$ particles of corresponding types $1,2,\ldots,k$. At any given time $t$, each pair of clusters with respective size vectors ${\bf x}$ and ${\bf y}$ will coalesce into a cluster of size ${\bf x+y}$ with rate $K({\bf x},{\bf y})/n$, where
(1) $K({\bf x},{\bf y})=\langle{\bf x}|V|{\bf y}\rangle.$
The last merger will create a cluster of size $\,{\boldsymbol{\alpha}}n$. Clusters coalesce independently of other clusters.
The kernel $K({\bf x},{\bf y})$ defined in (1) will be referred to as the vector-multiplicative kernel. The kernel is symmetric
$K({\bf x},{\bf y})=K({\bf y},{\bf x})\quad\text{ for all vectors }~{}{\bf x},{\bf y}$
and bilinear
$K(c_{1}{\bf x}+c_{2}{\bf y},{\bf z})=c_{1}\,K({\bf x},{\bf z})+c_{2}\,K({\bf y},{\bf z})\quad\text{ for all vectors }~{}{\bf x},{\bf y},{\bf z}~{}~{}\text{ and scalars }~{}c_{1},c_{2}.$
In the vector-multiplicative coalescent process, let $\zeta_{\bf x}^{[n]}(t)$ denote the number of clusters of size ${\bf x}$ at time $t\geq 0$. The initial values are $\zeta_{\bf x}^{[n]}(0)=\alpha_{i}n$ if ${\bf x}={\bf e}_{i}$, where ${\bf e}_{i}$ is the $i$-th standard basis vector and $\zeta_{\bf x}^{[n]}(0)=0$ otherwise. The process
${\bf ML}_{n}(t)=\Big{(}\zeta_{\bf x}^{[n]}(t)\Big{)}_{{\bf x}\in\mathbb{Z}_{+}^{k}:\langle{\bf x}|{\bf 1}\rangle>0}$
that counts clusters of all sizes in the vector-multiplicative coalescent process is the corresponding Marcus-Lushnikov process. We define $\zeta_{\bf x}(t)=\lim_{n\rightarrow\infty}\frac{\zeta_{\bf x}^{[n]}(t)}{n}$ to yield the deterministic limiting fraction of clusters of each size ${\bf x}$. The differential equations that govern the dynamics of $\zeta_{\bf x}(t)$ are called the Smoluchowski coagulation system of equations and are given by
(2) ${d\over dt}\zeta_{\bf x}(t)=-\zeta_{\bf x}\sum_{y}\zeta_{\bf y}\langle{\bf x}|V|{\bf y}\rangle+\frac{1}{2}\sum_{{\bf y},{\bf z}\,:{\bf y}+{\bf z}={\bf x}}\langle{\bf y}|V|{\bf z}\rangle\zeta_{{\bf y}}\zeta_{{\bf z}}$
with initial conditions $\zeta_{\bf x}(0)=\alpha_{i}$ if $x={\bf e}_{i}$, and $\zeta_{\bf x}(0)=0$ otherwise. These equations describe the evolution of the size distribution of clusters in a system where clusters merge multiplicatively as defined in (1). More specifically, the first term on the right hand side in (2) corresponds to the depletion of clusters of size x after merging with other clusters and the second term corresponds to coagulation of smaller clusters to produce one of size x. For further details on Smoluchowski equations in various settings including applications, we refer the reader to [21, 18, 10, 19, 5, 14, 7].
1.2. The solution to the Smoluchowski equations
In [16], it was shown that the solutions to the Smoluchowski equations and the following modified Smoluchowski equations (MSE) given by
(3) ${d\over dt}\zeta_{\bf x}(t)=-\zeta_{\bf x}\langle{\bf x}|V|\boldsymbol{\alpha}\rangle+\frac{1}{2}\sum_{{\bf y},{\bf z}\,:{\bf y}+{\bf z}={\bf x}}\langle{\bf y}|V|{\bf z}\rangle\zeta_{{\bf y}}\zeta_{{\bf z}}$
with the same initial conditions coincide as long as the second order moments are convergent. The unique solution $\zeta_{\bf x}(t)$ of the above modified Smoluchowski equations was derived in [16]. Specifically, for a vector ${\bf x}\in\mathbb{Z}_{+}^{k}$ let $\,{\bf x}!=x_{1}!x_{2}!\ldots x_{k}!\,$ and for vectors ${\bf a}$ and ${\bf b}$ in $\mathbb{R}^{k}$ let $\,{\bf a}^{\bf b}=a_{1}^{b_{1}}a_{2}^{b_{2}}\ldots a_{k}^{b_{k}}\,$ whenever $a_{i}^{b_{i}}$ is uniquely defined for all $i$. Now, consider a complete graph $K_{k}$ consisting of vertices $\{1,\ldots,k\}$ with weights $w_{i,j}=w_{j,i}\geq 0$ assigned to its edges $[i,j]$ ($i\not=j$). Let the weight $W(\mathcal{T})$ of a spanning tree $\mathcal{T}$ be the product of the weights of all of its edges. Finally, let $\tau(K_{k},w_{i,j})=\sum\limits_{\mathcal{T}}W(\mathcal{T})$ denote the weighted spanning tree enumerator, i.e., the sum of weights of all spanning trees in $K_{k}$. In the case $w_{i,j}\equiv 1$, the weighted spanning tree enumerator simply equals the number of spanning trees of the graph. In terms of these notations, the solution to the modified Smoluchowski equations that yield the limiting fraction of clusters of size ${\bf x}$ is given by
(4) $\zeta_{\bf x}(t)={1\over{\bf x}!}\boldsymbol{\alpha}^{\bf x}{\tau(K_{k},x_{i}x_{j}v_{i,j})\over{\bf x}^{\bf 1}}(V{\bf x})^{{\bf x}-{\bf 1}}e^{-\langle{\bf x}|V|\boldsymbol{\alpha}\rangle t}t^{\langle{\bf x}|\boldsymbol{1}\rangle-1}.$
Outline of the paper.  The rest of the paper is organized as follows: Section 2 is devoted to a short proof of equation (4) based on the theory of random graphs, which could be of independent interest, as it provides a probabilistic perspective on the Smoluchowski equations. In Section 3, we discuss gelation and its connection to the multidimensional Lambert-Euler inversion. In Section 4, we recall some basic concepts from the theory of multi-type branching processes. We conclude by stating the connection between the vector-multiplicative coalescent process and multi-type Poisson branching processes and its applications.
2. Random graphs and an alternative proof of equation (4)
The following alternative proof uses the connection between the vector multiplicative coalescent process $C$ and the theory of random graphs. This connection works as follows: clusters of particles in $C$ correspond to small connected components of a certain random graph $G$. The so-called giant component of $G$ corresponds to the gel in $C$ for details, see the next section. The correspondence between $C$ and $G$ was first described in [16], to which we refer the reader for a full account. Here, we include only enough detail to describe the proof of (4).
The random graph $G$ models the coalescent process at a fixed time $t$. The $n$ particles in $C$ correspond to the $n$ vertices of $G$, and, as before, we will let $n\to\infty$. The numbers $\alpha_{i}n$ of particles of each type in $C$ correspond to the sizes of the parts $A_{i}$ of $G$, so that $G$ consists of (almost exactly) $n$ vertices, divided into $k$ parts $A_{1},\ldots,A_{k}$, in such a way that part $A_{i}$ contains $\alpha_{i}n$ vertices. Each potential edge $x_{i}x_{j}$ of $G$ (with $x_{i}\in A_{i}$ and $x_{j}\in A_{j}$) is included in $E(G)$, the edge set of $G$, with probability
$1-e^{-v_{ij}t/n}=\frac{v_{ij}t}{n}+O\left(\left(\frac{v_{ij}t}{n}\right)^{2}\right).$
Here, $V=(v_{ij})$ is the partition interaction strength matrix defined above. The exponential formulation is necessary so that $G$ can be seen as a snapshot of a certain memoryless process. Finally, for our purposes, $t$ will be a constant, not depending on $n$. Note that if $k=\alpha_{1}=v_{11}=1$ then ${\bf G}$ is (modulo the approximation above) just the Erdős-Rényi graph $G(n,t/n)$, with average degree (approximately) $t$. (Also note that we are allowing edges inside the parts, so that our graphs are not necessarily $k$-partite.)
With $\alpha$ and $V$ fixed, there will be a critical value $t^{\prime}$ of $t$ below which there is almost surely no giant component, and above which there is almost surely a giant component. Denote by $D[\alpha_{i}]$ the diagonal matrix with the $\alpha_{i}$ on the diagonal. Then, this critical value $t^{\prime}$ is such that the largest eigenvalue of the matrix $t^{\prime}VD[\alpha_{i}]$ equals 1 (i.e., $t^{\prime}$ is the reciprocal of the spectral radius of $VD[\alpha_{i}]$).
Next, again with $V$ and $\alpha$ fixed, and for given values of ${\bf x}=(x_{1},\ldots,x_{k})$ and $t$, we seek the scaled number of finite components in ${\bf G}$ containing $x_{i}$ vertices of type $i$. By “scaled”, we mean the number of such components divided by $n$. Call (the limit as $n\to\infty$ of) this scaled quantity $\phi_{\bf x}(t)$. Note that the process undergoes a qualitative change when $t=t^{\prime}$, but, for a given ${\bf x}$, the functions $\phi_{\bf x}(t)$ are in fact continuous (and probably even differentiable) for all $t$.
The following theorem, connecting the process $C$ and the graph $G$, appears in [16] (see also [1]).
Theorem 2.1.
[1] For given ${\boldsymbol{\alpha}},V$ and for any ${\bf x}$ and $t<t^{\prime}=T_{\rm gel}$, we have
$\phi_{\bf x}(t)=\zeta_{\bf x}(t),$
where the functions $\zeta_{\bf x}(t)$ are the solutions of the modified Smoluchowski equations (3) above.
Consequently, at least in the pre-gelation interval, we can use the random graph formulation to compute $\zeta_{\bf x}(t)$. We choose a method described on page 188 in Chapter 11 of [2].
For simplicity, we first address the case where $k=\alpha_{1}=v_{11}=1$. In this case, setting ${\bf x}=l$, we have
(5) $\zeta_{l}(t)=\phi_{l}(t)=\frac{l^{l-2}t^{l-1}e^{-lt}}{l!}$
The random graphs proof of this is as follows. We calculate $p_{l}(t)=l\phi_{l}(t)$, which is the probability that a given vertex lies in a connected component $G_{l}$ of size $l$ (which, with probability tending to 1, is a tree - see [6] for a proof). Now, setting $p=t/n$ for the edge probability, we see that:
  • There are $\binom{n-1}{l-1}$ choices for the other vertices in $G_{l}$
  • There are $l^{l-2}$ labeled trees on these chosen vertices
  • Each tree is present with probability $p^{l-1}$
  • The edges between $G_{l}$ and the other vertices are missing with probability $(1-p)^{l(n-l)}$
Therefore, as $n\to\infty$,
$p_{l}(t)$ $\sim\binom{n-1}{l-1}\cdot l^{l-2}\cdot p^{l-1}\cdot(1-p)^{l(n-l)}\sim\frac{n^{l-1}}{(l-1)!}\cdot l^{l-2}\cdot\left(\frac{t}{n}\right)^{l-1}\cdot e^{-pln}$
$=\frac{1}{(l-1)!}\cdot l^{l-2}\cdot t^{l-1}\cdot e^{-tl}=\frac{l^{l-1}t^{l-1}e^{-lt}}{l!},$
which, recalling that $p_{l}(t)=l\phi_{l}(t)$, establishes (5).
Next we consider the general case. For this we once again require the weighted spanning tree enumerator defined above, although here it arises in a slightly different form. Write $K_{\bf x}(V)$ for the complete graph on $\langle{\bf x}|\boldsymbol{1}\rangle$ vertices with $x_{i}$ vertices of type $i$, and where the weight of an edge between vertices of types $i$ and $j$ is $v_{ij}$. Suppressing the dependence on $V$, write $T_{\bf x}$ for the sum of the weights of all the spanning trees of $K_{\bf x}(V)$. Then, from Theorem 3.8 in [16],
$T_{\bf x}=\frac{\tau(K_{k},x_{i}x_{j}v_{i,j})}{{\bf x}^{\bf 1}}(V{\bf x})^{{\bf x}-{\bf 1}}.$
Now, generalizing the argument above, pick a random vertex $v$. The probability that $v$ is of type $i$ is $\alpha_{i}$. For simplicity of notation, suppose that $i=1$. We aim to compute $p_{\bf x}(t)$, which is the probability that $v$ lies in a connected component $G_{\bf x}$ of type ${\bf x}$. Again, with probability tending to 1, $G_{\bf x}$ is a tree [6]. Omitting floor and ceiling functions:
  • There are $\binom{\alpha_{1}n}{x_{1}-1}\binom{\alpha_{2}n}{x_{2}}\cdots\binom{\alpha_{k}n}{x_{k}}$ choices for the other vertices in $G_{\bf x}$
  • These form a tree with probability $T_{\bf x}\cdot(t/n)^{\langle{\bf x}|\boldsymbol{1}\rangle-1}$
  • The edges from $G_{\bf x}$ to $G\setminus G_{\bf x}$ are missing with probability $\sim\exp(-\sum_{i}\sum_{j}x_{i}v_{ij}\alpha_{j}t)$
Therefore, as $n\to\infty$,
$p_{\bf x}(t)$ $\sim\langle{\bf x}|\boldsymbol{1}\rangle\frac{n^{\langle{\bf x}|\boldsymbol{1}\rangle-1}\alpha_{1}^{x_{1}}\cdots\alpha_{k}^{x_{k}}}{x_{1}!\cdots x_{k}!}\cdot T_{\bf x}\cdot\left(\frac{t}{n}\right)^{\langle{\bf x}|\boldsymbol{1}\rangle-1}\cdot{\rm exp}\left(-\sum_{i}\sum_{j}x_{i}v_{ij}\alpha_{j}t\right)$
$=\langle{\bf x}|\boldsymbol{1}\rangle\frac{\alpha_{1}^{x_{1}}\cdots\alpha_{k}^{x_{k}}}{x_{1}!\cdots x_{k}!}\cdot T_{\bf x}\cdot t^{\langle{\bf x}|\boldsymbol{1}\rangle-1}\cdot{\rm exp}\left(-\sum_{i}\sum_{j}x_{i}v_{ij}\alpha_{j}t\right),$
so that, since
$p_{\bf x}(t)=\phi_{\bf x}(t)\langle{\bf x}|\boldsymbol{1}\rangle,$
we have
$\zeta_{\bf x}(t)=\phi_{\bf x}(t)=\frac{\alpha_{1}^{x_{1}}\cdots\alpha_{k}^{x_{k}}}{x_{1}!\cdots x_{k}!}\cdot T_{\bf x}\cdot t^{\langle{\bf x}|\boldsymbol{1}\rangle-1}\cdot{\rm exp}\left(-\sum_{i}\sum_{j}x_{i}v_{ij}\alpha_{j}t\right),$
proving (4).
3. Gelation and Lamber-Euler Inversion
It is natural to ask if the total mass for a system described by the Smoluchowski coagulation model is conserved. In [13, 20, 3, 22, 23], it was shown that for certain classes of coagulation kernels, this is not always possible as a system can undergo a phase transition. For example, Hendriks et al. [11] studied this phase transition for the model with the multiplicative kernel (in one-dimensional setting), and they showed that the total mass decreases after a finite time, observing a surprising phenomenon known as gelation. Broadly speaking, gelation is the break down of conservation of mass in the system and can be physically interpreted as being caused by the formation of a giant cluster called the gel.
For the coalescent process defined in section 1.1, the initial total mass vector is assumed to be $\,\sum_{\bf x}\zeta_{\bf x}(0)|{\bf x}\rangle=|\boldsymbol{\alpha}\rangle$. Therefore, we define the gelation as the loss of total mass $\,\sum_{\bf x}\zeta_{\bf x}(t)|{\bf x}\rangle\,$ of the system after a critical time $T_{gel}$ called gelation time. That is
(6) $T_{gel}=\inf\Big{\{}t>0\,:\,\sum_{\bf x}\zeta_{\bf x}(t)|{\bf x}\rangle<|\boldsymbol{\alpha}\rangle\Big{\}}.$
In [16], the existence of gelation was proved for the vector multiplicative coalescent process and the gelation time was given by
(7) $T_{gel}={1\over\rho(VD[\alpha_{i}])},$
where $\rho(A)$ denotes the spectral radius of matrix $A$.
The existence of gelation for the vector multiplicative coalescent process was proved by solving a multidimensional Lambert-Euler inversion problem. This is a higher dimensional generalization of the equation originally studied by Lambert and followed up by Euler that gave rise to the well known Lambert W function [17, 9]. For a given vector $\boldsymbol{\alpha}\in(0,\infty)^{k}$, consider the regions
(8) $R_{0}=\left\{\boldsymbol{\alpha}\in(0,\infty)^{k}:\,\rho\left(VD[\alpha_{i}]\right)<1\right\}$
its closure within $(0,\infty)^{k}$,
(9) $\overline{R}_{0}=\left\{\boldsymbol{\alpha}\in(0,\infty)^{k}:\,\rho\left(VD[\alpha_{i}]\right)\leq 1\right\},$
and the complement of $\overline{R}_{0}$ within $(0,\infty)^{k}$,
(10) $R_{1}=\left\{\boldsymbol{\alpha}\in(0,\infty)^{k}:\,\rho\left(VD[\alpha_{i}]\right)>1\right\}.$
The following theorem implies the existence of a unique solution to the multidimensional Lambert-Euler problem (see [16, Theorem 1.1]).
Theorem 3.1 (Multidimensional Lambert-Euler inversion).
Consider a nonnegative irreducible symmetric matrix $V\in\mathbb{R}^{k\times k}$. Then, for any given $\boldsymbol{\alpha}\in(0,\infty)^{k}$ and $t\geq 0$, there exists a unique vector ${\bf y}\in\overline{R}_{0}$ such that
(11) $y_{i}e^{-\langle{\bf e}_{i}|V|{\bf y}\rangle}=\alpha_{i}te^{-\langle{\bf e}_{i}|V|\boldsymbol{\alpha}t\rangle}\qquad i=1,\ldots,k.$
Moreover, if $\boldsymbol{\alpha}t\in\overline{R}_{0}$, then ${\bf y}=\boldsymbol{\alpha}t$. If $\boldsymbol{\alpha}t\in R_{1}$, then ${\bf y}<\boldsymbol{\alpha}t$ ($y_{i}<\alpha_{i}t$ for all $i$), i.e., ${\bf y}$ is the smallest solution of (11).
Finally, in [16], it was shown that the total mass $\sum_{\bf x}\zeta_{\bf x}(t)|{\bf x}\rangle$ of the vector multiplicative coalescent process could be expressed in terms of the solution of the multidimensional Lambert-Euler inversion problem as follows:
(12) $\sum_{\bf x}\zeta_{\bf x}(t)|{\bf x}\rangle={1\over t}|{\bf y}(t)\rangle$
Then, Theorem 3.1 proves the existence of the gelation phenomenon for the vector multiplicative coalescent process and the gelation time given in (7). Specifically,
$\sum_{\bf x}\zeta_{\bf x}(t)|{\bf x}\rangle\left\{\begin{array}[]{ll}=|{\boldsymbol{\alpha}}\rangle&\mbox{if $t\leq 1/\rho\left(VD[\alpha_{i}]\right)$}\\ <|{\boldsymbol{\alpha}}\rangle&\mbox{if $t>1/\rho\left(VD[\alpha_{i}]\right)$}\end{array}\right.$
Remark 3.2.
The one dimensional Lambert-Euler inversion reads as
$x(t)=\min\{x\geq 0\,:\,xe^{-x}=te^{-t}\}.$
In their now famous paper [8], Erdős and Rényi used the function $x(t)$ to establish the formation of a giant component in the theory of random graphs. In fact, in random graph setting, the quantity $x(t)$ can be interpreted as the the average degree outside of the giant component in an Erdős-Rényi graph. It would be interesting to see if (11) can be used to obtain refined information about the proportion of vertices in the giant component of a random multipartite graphs, in the same spirit as the classical case. Note that in this one-dimensional setting, $x(t)=t$ on $(0,1]$, and $x(t)<t$ on $(1,\infty)$. Moreover, $x(t)\downarrow 0$ monotonically as $t\to\infty$. On the other hand, the post-gelation behavior of the multi-dimensional solution to (11) can differ from the one-dimensional case. For example, consider the bipartite case with $V=\begin{vmatrix}0&1\\ 1&0\end{vmatrix}$ and ${\boldsymbol{\alpha}}=(15,2)$. Then, the critical time is given by $T_{gel}=\frac{1}{\sqrt{30}}$, and for $t>T_{gel}$, the graph of $y(t)=(y_{1}(t),y_{2}(t))$ is as follows.
Figure 1. Post-gelation behavior of Lambert-Euler inversion
Observe that the function $y_{1}(t)$ decreases, increases, and finally decreases as $t$ increases. Numerical simulation seems to imply that there is a critical value for the ratio of the parameters in ${\boldsymbol{\alpha}}$ after which the solutions $(y_{1}(t),y_{2}(t))$ are no longer monotonically decreasing.
4. Multi-Type Poisson Branching Processes
In this and the next sections, we give an overview of the relevant theory of multi-type branching processes which is a generalization of the classical Galton-Watson process, extending it to multiple types of individuals. A full theory can be found in [4]. This multi-type model is useful in various fields, including biology, physics, and epidemiology, to study populations with different types of individuals and how these types evolve over time. Recall that in the classical Galton-Watson process, there is only one type of individual, whereas in the multi-type case, there are $k$ different types: $1,2,\ldots,k$. We shall assume that the process starts with only one individual of a specific type initially, i.e. ${\bf Z}_{0}={\bf e}_{i}$. Let ${\bf Z}_{n}=(Z_{1n},Z_{2n},\ldots Z_{kn})$ be the $k$-dimensional vector representing the number of individuals of each type in the $n$th generation of the branching process. The number of each type are independent random variables. Let ${\bf z}=(z_{1},\ldots,z_{k})$. For each type $i=1,2,\ldots,k$, denote by $p_{i}(\bf z)$ the probability that an individual of type $i$ has the offspring vector ${\bf z}$. Note that $\sum_{\bf z}p_{i}(\bf z)=1$. These probabilities define the offspring distribution for individuals of type $i$.
To analyze the probabilistic structure of the offspring distribution, we can use generating functions. The probability generating functions for a multi-type branching process with the offspring distribution $(p_{i})$ is defined as
(13) $f_{i}(s_{1},\ldots,s_{k})=\sum_{{\bf z\in\mathbb{N}_{0}^{k}}}p_{i}({\bf z})s_{1}^{z_{1}}s_{2}^{z_{2}}\cdots s_{k}^{z_{k}},\,\,\,\,\,\,\,\,\,\,|s_{1}|,\ldots,|s_{k}|\leq 1.$
We say $f_{i}$ is the probability generating function for the offspring distribution of individuals of type $i$.
For a $k$-type branching process, define the matrix of means as the $k\times k$ matrix ${\bf M}=[m_{ij}]$ such that
(14) $m_{ij}=\mathbb{E}[Z_{j1}\,|\,{\bf Z}_{0}={\bf e}_{i}].$
In other words, the entry $m_{ij}$ represents the average number of offspring of type $j$ produced by an individual of type $i$ in one generation. In the case where the entries $m_{i,j}$ vary, we have a multi-type inhomogenous branching process.
Let us now consider the multi-type Poisson branching process, a special case of the multi-type branching process where the number of offspring of each type follows a Poisson distribution. Then the entries of the corresponding mean matrix $M$ are the parameters of the Poisson distributions, i.e. $Z_{ij}\sim\text{Poi}(m_{ij})$. This process has offspring distributions of the form
(15) $p_{i}({\bf z})=\prod_{j=1}^{k}e^{-m_{ij}}\frac{(m_{ij})^{z_{j}}}{z_{j}!}.$
Applying (15) to (13) and the independence of number of offsprings of each type yields the following form for probability generating function of the multi-type Poisson branching process.
(16) $f_{i}({\bf s})=e^{\langle{\bf e}_{i}\,|{\bf M}\,|{\bf s}-\boldsymbol{1}\rangle}.$
4.1. Extinction Probability
As gelation is to vector multiplicative coalescent processes, a fundamental question in the study of branching processes is whether the process becomes extinct as the number of generations $n$ tends to infinity. Recall that in the classical case, if the expected number of the offspring distribution is less than (or equal to) 1, then eventually the population dies out. But if the expected number is greater than 1, then the probability of extinction is the smallest non-negative solution to the equation $s=\varphi(s)$, where $\varphi$ is the probability generating function of the corresponding offspring distribution.
For a multi-type branching process with offspring distribution $({\bf Z}_{n})$, let ${\boldsymbol{\eta}}=(\eta_{1},\ldots,\eta_{k})$ denote the extinction probability vector, where $\eta_{i}$ denote the extinction probability of an individual of type $i$ when the process begins with ${\bf Z}_{0}={\bf e}_{i}$. In this case, the phase transition for ${\boldsymbol{\eta}}$ is related to the spectral radius of the mean matrix ${\bf M}$, as stated in the following theorem (see [4]).
Theorem 4.1.
Let $f_{i}({\bf s})$ be the probability generating function and ${\bf M}$ be the matrix of means of a positive regular and non-singular multi-type branching process. Let $\rho({\bf M})$ denote the spectral radius of ${\bf M}$. Furthermore, let ${\boldsymbol{\eta}}$ be the vector of extinction probabilities.
(a) If $\rho({\bf M})\leq 1$, then ${\boldsymbol{\eta}}={\bf 1}$, the $k$-dimensional vector $(1,1,\ldots,1)$.
(b) If $\rho({\bf M})>1$, then ${\boldsymbol{\eta}}$ is the smallest (by component) solution less than ${\bf 1}$ of the fixed point equations
(17) $f_{i}({\bf s})=s_{i}\,,\,\,\forall\,i=1,2,\ldots,k.$
In the case of the multi-type Poisson branching process, the fixed point equations (17) have the form
(18) $e^{\langle{\bf e}_{i}\,|{\bf M}\,|{\bf s}-\boldsymbol{1}\rangle}=s_{i}$
which we will refer to as the Poisson fixed point equations.
5. Correspondence between Gelation and Extinction Probability with Applications
The bridge between the phase transitions of these two stochastic models stems from the equivalence of the multidimensional Lambert-Euler inversion equations (11) and the Poisson fixed point equations (18) with ${\bf M}=VD[\alpha_{i}]t$. Specifically,
(19) $e^{\langle{\bf e}_{i}\,|VD[\alpha_{i}]t\,|{\bf s}-\boldsymbol{1}\rangle}=s_{i}\hskip 21.68121pt\Longleftrightarrow\hskip 21.68121pty_{i}e^{-\langle{\bf e}_{i}|V|{\bf y}\rangle}=\alpha_{i}te^{-\langle{\bf e}_{i}|V|\boldsymbol{\alpha}t\rangle}$
when $s_{i}=y_{i}/(\alpha_{i}t)$. The identification of the matrix of means ${\bf M}$ in the multi-type Poisson branching process in terms of the time parameter $t$ of the vector multiplicative coalescent process implies that $t$ is the parameter for a continuous sequence of branching processes rather than the usual time (generations) view of branching processes.
Using this equivalence we get our first application that yields the gelation phenomenon for the vector multiplicative coalescent process discussed in Section 3. We state the result again in the theorem below followed by the short new proof that follows immediately from this equivalence.
Theorem 5.1.
Let $\zeta_{\bf x}(t)$ be the solution to the modified Smoluchowski equations which represents the limiting fraction of clusters of size ${\bf x}$ in the vector multiplicative coalescent process.
(a) If $\rho(VD[\alpha_{i}]t)\leq 1$, then the total mass $\sum_{\bf x}\zeta_{\bf x}(t)|{\bf x}\rangle=|{\boldsymbol{\alpha}}\rangle$.
(b) If $\rho(VD[\alpha_{i}]t)>1$, then the total mass $\sum_{\bf x}\zeta_{\bf x}(t)|{\bf x}\rangle<|{\boldsymbol{\alpha}}\rangle$.
Consequently, the gelation time for the vector multiplicative coalescent process is given by $T_{gel}={1\over\rho(VD[\alpha_{i}])}$.
Proof.
Applying the identification ${\bf M}=VD[\alpha_{i}]t$ to Theorem 4.1 yields
(a) If $\rho(VD[\alpha_{i}]t)\leq 1$, then ${\boldsymbol{\eta}}={\bf 1}$, the $k$-dimensional vector $(1,1,\ldots,1)$.
(b) If $\rho(VD[\alpha_{i}]t)>1$, then ${\boldsymbol{\eta}}$ is the smallest solution less than ${\bf 1}$ of the fixed point equations
$e^{\langle{\bf e}_{i}\,|VD[\alpha_{i}]t\,|{\bf s}-\boldsymbol{1}\rangle}=s_{i}$
Then the equivalence (19) implies that $\boldsymbol{\alpha}t\boldsymbol{\eta}={\bf y}$ which completes the proof. ∎
Remark 5.2.
In a recent work [12], Hoogendijk et al. described an alternative approach, which relies on establishing the equivalence between the Smoluchowski coagulation equations (see equation (2)) and the inviscid Burgers equation using branching process techniques.
The second application of the equivalence (19) is a new series expression for the extinction probabilities of the Poisson branching process. From the equivalence (19), we have the extinction probability $\eta_{\ell}=y_{\ell}/(\alpha_{\ell}t)$ where $y_{\ell}$ is the smallest solution to the Lambert-Euler inversion equations. Then by the total mass expression (12) and the explicit solution to the modified Smoluchowski equations (4), we have
(20) $\eta_{\ell}=\frac{1}{\alpha_{\ell}}\frac{y_{\ell}}{t}=\frac{1}{\alpha_{\ell}}\sum_{{\bf x}}\zeta_{{\bf x}}(t)x_{\ell}=\frac{1}{\alpha_{\ell}}\sum_{{\bf x}}{1\over{\bf x}!}\boldsymbol{\alpha}^{\bf x}{\tau(K_{k},x_{i}x_{j}v_{i,j})\over{\bf x}^{\bf 1}}(V{\bf x})^{{\bf x}-{\bf 1}}e^{-\langle{\bf x}|V|\boldsymbol{\alpha}\rangle t}t^{\langle{\bf x}|\boldsymbol{1}\rangle-1}x_{\ell}.$
From the relation ${\bf M}=VD[\alpha_{i}]t$,
$\langle{\bf x}|V|\boldsymbol{\alpha}\rangle t=\langle{\bf x}|{\bf M}|\boldsymbol{1}\rangle$
and
$\boldsymbol{\alpha}^{\bf x}(V{\bf x})^{{\bf x}-{\bf 1}}=\frac{\boldsymbol{\alpha}^{\boldsymbol{1}}}{t^{\langle{\bf x}|\boldsymbol{1}\rangle-k}}({\bf M}{\bf x})^{{\bf x}-{\bf 1}}$
Moreover,
$\tau(K_{k},x_{i}x_{j}v_{i,j})=\frac{1}{t^{k-1}}\ \tau\left(K_{k},x_{i}x_{j}\frac{m_{i,j}}{\alpha_{j}}\right)$
Substituting the above three equations into (20) yields the following new result for the extinction probabilities of the multi-type Poisson branching process.
Theorem 5.3.
Suppose the matrix of means ${\bf M}$ for a multi-type Poisson branching process can be expressed as ${\bf M}=VD[\alpha_{i}]t$ where $V$ is a nonnegative, irreducible and symmetric matrix, ${\boldsymbol{\alpha}}=(\alpha_{1},\alpha_{2},\ldots,\alpha_{k})\in(0,\infty)^{k}$, and $t$ is a nonnegative real number. Then the extinction probabilities of this process ${\boldsymbol{\eta}}=(\eta_{1},\ldots,\eta_{k})$ have the series form
$\eta_{\ell}=\frac{1}{\alpha_{\ell}}\sum_{{\bf x}}{\boldsymbol{\alpha}^{\boldsymbol{1}}\over{\bf x}!\ {\bf x}^{\bf 1}}\ \tau\left(K_{k},x_{i}x_{j}\frac{m_{i,j}}{\alpha_{j}}\right)e^{-\langle{\bf x}|{\bf M}|\boldsymbol{1}\rangle}x_{\ell}\,,\,\,\,\,\,\,\forall\ell=1,2\ldots,k.$
In particular, the homogenous multi-type Poisson branching process where $\boldsymbol{\alpha}={\bf 1}$ and $V$ is matrix with all entries $1$, the extinction probabilities have the form
(21) $\eta_{\ell}=\sum_{{\bf x}}\frac{1}{{\bf x}!\ {\bf x}^{\bf 1}}\ \tau\left(K_{k},x_{i}x_{j}t\right)e^{-\langle{\bf x}|\boldsymbol{1}\rangle kt}x_{\ell}\,,\,\,\,\,\,\,\forall\ell=1,2\ldots,k.$